OpenStudy (anonymous):

On a fair spinner,the probability of spinning blue 1/3.The probability of spinning blue or green is 1/2.What is probability of spinning green?

OpenStudy (sandra):

this is a mutually exclusive probability problem

OpenStudy (sandra):

i.e. it's impossible for both events to happen at the same time

OpenStudy (anonymous):

yeea i know

OpenStudy (sandra):

and hence we know that P(A or B) = P(A) + P(B)

OpenStudy (sandra):

so P(B or G) = P(B) + P(G) 1/2 = 1/3 + P(G) 1/2 - 1/3 = P(G) 3/6 - 2/6 = 1/6 = P(G)

OpenStudy (sandra):

so probability of green is 1/6

OpenStudy (anonymous):

oh i have more quesitons

OpenStudy (sandra):

I'd say ask them separately? that way in case I can't answer it people won't think it's already been answered.

OpenStudy (anonymous):

oh